# The Mittag-Leffler theorem for proper minimal surfaces and directed meromorphic curves

### v1

@@ 6.  The Mittag-Leffler theorem for meromorphic $A$-immersions
(7.2) $\mu_{j}:=\min\left\{||F(p)||_{\infty}\colon p\in W_{j}\right\}-1\geq-1$
and note that, by (7.1), $\lim_{j\to\infty}\mu_{j}=+\infty$. As in the proof of Theorem 5.1, take a normal exhaustion $\{M_{j}\}_{j}$ of $M$ satisfying the same properties as (5.1), and assume that $W_{j}$ is the unique component of $U$ lying in $M_{j}\setminus M_{j-1}$ for every $j\geq 1$. Choose a full holomorphic $A$-embedding $F^{0}\colon M_{0}\to\mathbb{C}^{n}$ such that $F^{0}$ has no zeros on $bM_{0}$, let $0<\delta_{0}<\min\{||F^{0}(p)||_{\infty}\colon p\in bM_{0}\}$ and $\varepsilon_{0}:=\varepsilon/2$. We shall inductively construct a decreasing sequence of positive numbers $\{\varepsilon_{j}\}_{j}$, a sequence of positive numbers $\{\delta_{j}\}_{j}$ with $\lim_{j\to\infty}\delta_{j}=+\infty$, and a sequence of full meromorphic $A$-immersions $\{F^{j}\colon M_{j}\setminus(E\cap M_{j})\to\mathbb{C}^{n}\}_{j}$ of class $\mathscr{I}_{A}(M_{j}|E\cap M_{j})$, such that properties (1j)–(5j) in the proof of Theorem 5.1 as well as the following ones hold for all $j\geq 1$.
@@
    $||F^{j}(p)||_{\infty}>\min\{\delta_{j-1},\mu_{j}\}$ for all $p\in M_{j}\setminus\mathring{M}_{j-1}$.
@@
(7.3) $||F^{j-1}||_{\infty}>\min\left\{\delta_{j-1},\mu_{j}\right\}\quad\textrm{on }\Gamma_{j},$
@@
(7.4) $||\widetilde{F}^{j}||_{\infty}>\min\{\delta_{j-1},\mu_{j}\}\quad\textrm{on }\widetilde{M}_{j-1}\setminus\mathring{M}_{j-1};$
@@
$G:=\lim_{j\to\infty}F^{j}\colon M\setminus E\to\mathbb{C}^{n}$
@@
(7.5) $||G-F^{j}||_{M_{j}}<2\varepsilon_{j}\leq\varepsilon,\quad j\geq 0;$
@@
For assertion (i), apply Lemma 7.2 instead of Lemma 7.3 and omit condition (8j) in the induction process. By (7j) and (7.5), it follows that $||G(p)||_{\infty}>\delta_{j}-\varepsilon$ for all $p\in bM_{j}$, which goes to +$\infty$+ ∞ as $j\to+\infty$. Taking into account (5.1) and that $G$ has an effective pole at every point in $E$, this implies that $G:M\setminus E\to\mathbb{C}^{n}$ is an almost proper map. ∎
@@ Acknowledgements
Research is partially supported by the State Research Agency (AEI) via the grants no. PID2020-117868GB-I00 and PID2023-150727NB-I00, and the “Maria de Maeztu” Unit of Excellence IMAG, reference CEX2020-001105-M, funded by MCIN/AEI/10.13039/501100011033/, Spain.

### v2

@@ 6. The Mittag-Leffler theorem for meromorphic $A$-immersions
[…]
@@
    $F^{j}(p)||_{\infty}>\min\{\delta_{j-1},\mu_{j}\}$ for all $p\in M_{j}\setminus\mathring{M}_{j-1}$.
@@
(7.3) $F^{j-1}||_{\infty}>\min\left\{\delta_{j-1},\mu_{j}\right\}\quad\textrm{on }\Gamma_{j},$
@@
(7.4) $\widetilde{F}^{j}||_{\infty}>\min\{\delta_{j-1},\mu_{j}\}\quad\textrm{on }\widetilde{M}_{j-1}\setminus\mathring{M}_{j-1};$
@@
$G:=\lim_{j\to\infty}F^{j}\colon M\setminus E\longrightarrow\mathbb{C}^{n}$
@@
(7.5) $G-F^{j}||_{M_{j}}<2\varepsilon_{j}\leq\varepsilon,\quad j\geq 0;$
@@
For assertion (i), apply Lemma 7.2 instead of Lemma 7.3 and omit condition (8j) in the induction process. By (7j) and (7.5), it follows that $G(p)||_{\infty}>\delta_{j}-\varepsilon$ for all $p\in bM_{j}$, which goes to $\infty$ as $j\to+\infty$. Taking into account (5.1) and that $G$ has an effective pole at every point in $E$, this implies that $G:M\setminus E\to\mathbb{C}^{n}$ is an almost proper map. ∎
@@ Acknowledgements
Research partially supported by the State Research Agency (AEI) via the grants no. PID2020-117868GB-I00 and PID2023-150727NB-I00, and the “Maria de Maeztu” Unit of Excellence IMAG, reference CEX2020-001105-M, funded by MICIU/AEI/10.13039/501100011033 and ERDF/EU, Spain.
We are grateful to two anonymous referees for their helpful remarks and suggestions which led to improve the exposition.